OpenStudy (anonymous):

find all of the solutions. system: {y = -1/2x^2 y = 3/4x-5

OpenStudy (p0sitr0n):

you can subtract the two equations here, so you dont have to isolate anything 0=1/2x^2+3/4x-5 now apply the quadratic formula a=1/2, b=3/4, c=-5 after solving, you get x=-4 or x =5/2 now plug those in the last equation (its simpler since no squaring) y=-3-5=-8 or y=15/8-5=-25/8 the solution space is then (-4,-8),(5/2,-25/8)

OpenStudy (anonymous):

So all of the solutions are (-4,-8), (5/2, -25/8)? How would you graph that on an axis? :) Thank you!

OpenStudy (anonymous):

@P0sitr0n

OpenStudy (p0sitr0n):

as two points